Circular No. 6315

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

GRO J1744–28 D. E. Vanden Berk, S. A. Severson, D. M. Cole, D. Q. Lamb, M. C.

Miller, R. C. Nichol, J. M. Quashnock, E. Bergeron, K. Gloria, and D. Long, University of Chicago, report follow-up analyses related to their discovery (IAUC 6310) of the optical/near-infrared counterpart of the variable radio source (IAUC 6307) in the XTE error box of GRO J1744–28: “During Feb. 5.534–5.582 UT, we carried out optical observations of a region of the sky that included the entire XTE error box (IAUC 6309) for GRO J1744–28, using the Astrophysical Research Consortium’s 3.5-m telescope at Apache Point Observatory. A re-analysis of our data (using our own photometric standards) shows that we reach limiting magnitudes g = 19.7 ± 0.3 and r = 20.5 ± 0.3. We see no new object or any significant brightening of any known object to these limits, compared to the COSMOS/NRL list and the European Southern Observatory (ESO) R copy of the U.K. Schmidt photographic plate. We now measure r = 20.1 ± 0.2 for the object reported by us on IAUC 6310 (we call this ‘star a’). From an ESO 2.2-m IRAC2 K′ image kindly provided by K. Glazebrook, C. Lidman, and P. Blanco, and using the preliminary photometric calibration provided by them, we measure K′ = 13.1 ± 0.3 for the source originally reported by us on IAUC 6310 (we call this ‘star A’). This is consistent with the value we reported earlier. Our astrometric solution is based on four stars in the COSMOS/NRL list that we are able to identify in all three passbands (g, r, and K′). At the position of the radio source, we detect nothing to the above limits in g and only one source (within 1′′) in r, that being at α = 17^{h}44^{m}36^{s}.94, δ = −28^{o}45′38′′.2 (equinox 2000.0) in the COSMOS/NRL reference frame (star a). In the ESO IRAC2 K′ image, we find four stars within 3′′ of the position of the radio source and list them here in increasing distance from that position (with α and δ position end figures): star A, 36^{s}.91, 37′′.8; star B, 36^{s}.93, 35′′.9; star C, 36^{s}.72, 38′′.5; star D, 36^{s}.87, 34′′.3.”

COMET C/1996 B2 (HYAKUTAKE) Total visual magnitude and coma diameter estimates: Feb. 12.18 UT,

8.6, – (J. M. Trigo, Castellon, Spain, 7×50 binoculars); 13.43, 8.6, 2′.0 (J. Bortle, Stormville, NY, 0.41-m reflector); 13.70, 7.9, 6′ (T. Lovejoy, Jimboomba, Queensland, 15×80 binoculars); 14.45, 8.3, – (A. Hale, Cloudcroft, NM, 10×50 binoculars); 14.45, 7.7, 8′ (C. S. Morris, Pine Mountain Club, CA, 20×80 binoculars).

1996 February 16  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1996 CBAT  Daniel W. E. Green